OpenStudy (anonymous):

How do you write the standard number 0.000041 in scientific notation?

OpenStudy (anonymous):

example: \(4000=4\cdot 10^3\) and \(0.003=3\cdot 10^{-3}\)

OpenStudy (tansheet):